OpenStudy (anonymous):

A root of x2 – 5x – 1 = 0 is

OpenStudy (anonymous):

Can be solved by completing the square. x^2 - 5x - 1 = 0 x^2 - 5x = 1 x^2 - 5x + (-5/2)^2 = 1 + (-5/2)^2 (x - 5/2)^2 = 29/4 sqrt[(x - 5/2)^2] = sqrt (29/4) x - 5/2 = [sqrt(29)]/2 x = 5/2 +- [sqrt(29)]/2 x = [5 +- sqrt (29)]/2

OpenStudy (midhun.madhu1987):